# From Label Error Detection to Correction: A Modular Framework and Benchmark for Object Detection Datasets

### v1

@@ IV-D Label Error Detection
MetaDetect. We use the MetaDetect meta classification model, which applies gradient boosting. We use the default values for all parameters and make predictions using -fold cross-validation on the respective hold-out set. The threshold, which determines whether predictions are considered to be false positives (FP), is set to 0.5. This threshold converts the IoU values of the predictions into binary labels that are used for training the meta classification model. For each predicted bounding box, the meta classification model predicts a probability for the prediction being correct, based on which we identify label errors. Metaclassification results are listed in the appendix.
@@
Instance-wise Loss Method. This label error detection method is based on object detection loss at the instance level. It requires a two-stage object detector, which is why we only apply it to Cascade R-CNN predictions. Based on images and the corresponding ground truth annotations, this method generates label error proposal boxes with a corresponding score and class prediction. We only consider proposals with a score of at least 0.01 for the class pedestrian.
@@
ObjectLab. The detected bounding boxes can be evaluated with respect to the overlooked score with ObjectLab for object detection. Originally, only confident predictions with a probability greater than 0.95 were considered, but to infer a score for every detected bounding box, this threshold is removed. All other parameters are set to their default values. Additionally, we consider the poor location scores of the original ground truth annotations. As there is only a single class under consideration, the swapped class label error does not occur here. As label quality scores are normalized and lower values indicate poorer label quality, we subtract these scores from 1 and interpret the result as a probability of the corresponding label issue occurring. The detected bounding boxes, as well as the original GT boxes, are then reviewed in order of probabilities.
@@
| Arbitrary | 611 | 176 | 231 | 113 |
| pixels | 398 | 159 | 183 | 110 |
| pixels | 131 | 115 | 63 | 87 |
IV-E Evaluation Metrics
Label Errors / FN in Original GT: We match predicted bounding boxes with the original ground truth (GT) using an threshold of 0.5, and with the validated GT (VGT) using a threshold of 0.1. These thresholds serve different purposes: the threshold is standard for checking agreement with original annotations, while the threshold allows for approximate matching against VGT, which acts as a proxy for reannotation. A prediction is considered a label error (i.e., a false negative in the original GT) if it has no match in the original GT but overlaps with the VGT. We classify errors into two types: completely missing () and misaligned boxes () see Table III. Matching is performed greedily by descending , and each VGT and GT annotation is matched at most once. We can similarly also analyze the FP in the original GT, for more details see the appendix.
Introduced Label Errors: To assess the risk of new errors introduced by labeling strategies, we compare their outcomes against the VGT. If a semantic validation strategy disagrees with the VGT on a bounding box’s label (e.g. one says pedestrian, the other not pedestrian), the box is misclassified. This may cause errors (FNs or FPs) to be missed or falsely introduced.
V Numerical Results
V-A Comparison of Original and Validated Annotations
Before comparing the VGT to the original ground truth, we first verify its quality. We evaluate three aspects: whether pedestrians were missed, whether duplicate bounding boxes exist, and how well the soft labels align with human perception. In a random sample of 200 images, we found that about 0.5% of pedestrians might have been missed, and about 0.8% of pedestrian bounding boxes were duplicates. See the appendix for the full analysis and visual examples. To evaluate the alignment of soft labels with human perception, we randomly sampled 200 bounding boxes from the VGT and let three expert annotators rate them. Each annotator could choose between the labels ’clearly’, ’maybe’, or ’no’ to indicate whether the bounding box contains a pedestrian. In Figure 4, these labels are compared with the corresponding soft label probability. The comparison shows that the categories match well with the expected scores. The category ’maybe’ has a mean and median of 0.48, which is almost perfectly calibrated, as we would expect a value around 0.5. For the category ’clearly’, we would expect a value around 0.8, which is consistent with the results. We conclude that the validated ground truth contains almost no missing pedestrians (about 0.5%), almost no duplicate bounding boxes (about 0.8%), and is well aligned with human perception.
We compare the validated annotations with the original annotations to determine the number of overlooked pedestrians and the number of misfitting or inaccurate bounding boxes. Overlooked pedestrians are VGT annotations that have an of zero with any original annotation. Misfitting boxes are VGT annotations for which there is an intersecting bounding box in the original ground truth, but with an below 0.5. We distinguish between large and small objects based on the minimum bounding box height, using thresholds of 25 or 40 pixels which relate to the difficulty levels defined in the KITTI benchmark. Additionally, we consider two different probability thresholds, and , for the soft labels in the VGT. The distribution of soft label probabilities is shown in 5(b) and illustrates that while most annotations show high agreement among annotators, some ambiguous cases occur. It is important to note that, unlike the original annotation protocol, the procedure described in subsection IV-B includes all pedestrians, regardless of their size or whether they fall into defined don’t care regions. As a result, we observe that the VGT also contains small objects that were not annotated in the original data, as shown in 5(a).
To ensure a fair comparison, we define a default configuration for the annotations under consideration. We include all objects that have an below 0.5 with any don’t care region and a bounding box height of at least 25 pixels. For the VGT annotations, we apply a soft label probability threshold of 0.5 to decide whether to include them. The resulting VGT contains 1,567 annotations for the class pedestrian, which is 671 more than in the original ground truth. Using this configuration, we identified a total of 497 label errors in the subset of 1,497 KITTI training images. Out of these, 398 pedestrians were missing entirely and 159 were annotated with inaccurate bounding boxes. We estimate that the original ground truth misses between 63 and 683 pedestrians, depending on the selected conditions such as minimum object size and probability threshold. This corresponds to between 7 – 76% relative to the 896 original pedestrian annotations. If we treat the overlooked and misaligned annotations as false negatives (FNs) relative to the original GT, we identify 293 FNs under strict conditions (applying a high confidence threshold, excluding don’t care regions, and enforcing size constraints) or 862 FNs when these restrictions are relaxed. This corresponds to a false negative rate (FNs / [FNs + Original GT]) ranging from 24.6% to 49%, depending on the evaluation setting. We provide visual examples of the 63 large objects that were missed, are not covered by don’t care regions, and have a soft label probability of 0.8 or higher in the appendix. In addition, we show random samples of other detected label errors such as inaccurate bounding box annotations and overlooked objects that either have lower confidence scores in the VGT or represent smaller pedestrians.
@@ V-B Label Error Detection and Correction
In 6(a), we compare the number of false negatives (FNs) in the original ground truth (GT) detected by various label error detection methods and labeling strategies. As outlined in section III, we define label errors as FNs in the original GT. We derive two key insights from the figure: (1) error detection methods make labeling more cost-efficient, and (2) there remains substantial room for improvement.
Among the evaluated methods, only YOLOX+MetaDetect consistently outperforms its base model. Cascade R-CNN+MetaDetect underperforms compared to Cascade R-CNN alone. ObjectLab performs significantly worse, likely due to our adaptation of its method; the original version uses a strict confidence threshold of 0.95, which we relaxed for this analysis. At such high thresholds, models provide highly confident predictions, but they cover fewer objects, limiting error discovery. For instance, using only predictions above 0.95 yields very few detected FNs, even in strong models. This highlights that restricting to high-confidence predictions alone is too limiting, even when using optimal detection techniques. Most detection-based approaches including YOLOX, Cascade R-CNN, and their error detection variants are able to find 150–200 FNs at a cost similar to or lower than the cheapest human annotation strategy. However, the base models saturate at 186 (YOLOX) and 260 (Cascade R-CNN) FNs, while the validated GT contains 333 FNs. This indicates that up to 44% of missing labels remain undetected. The gap increases under less restrictive matching conditions (e.g. without minimum box size), where up to 66% of FNs go unnoticed (see Appendix for full graph). This limitation stems primarily from the object detectors themselves, which constrain the set of proposed boxes. Labeling strategies show similar trends. As summarized in Table I, strategies using keypoint-to-box annotation generate fewer boxes and thus recover fewer FNs. Additional differences can be attributed to variations in label quality across correction strategies (see subsection V-C for ablations). While detectors limit recall, our RECD correction pipeline allows for more precise validation of the boxes they do propose and thus improving soft labeling quality. In summary, RECD identifies hundreds of label errors in less time than it takes humans to even annotate the bounding boxes. Yet, our analysis also reveals that up to 66% of potentially recoverable FNs remain undetected, highlighting significant research opportunities for improving label error detection.
@@ V-C Trade-offs Between Labeling Strategies and Label Quality
We evaluate the benefits and limitations of various labeling strategies in Figure 6. As expected, strategies cluster based on their bounding box annotation method discussed in the previous section. This is intuitive, since the number and quality of available boxes heavily influence the ability to identify FNs. Strategies using higher-quality soft labels consistently find more FNs. However, in less restrictive settings, e.g. relaxing the minimum box size (see appendix for full graph), even low-cost strategies can sometimes perform best. The second plot shows a clearer trend: the number of newly introduced errors decreases with increasing label quality. The lowest-cost semantic strategy (”Is pedestrian?”) introduces several hundred more errors than other methods. In contrast, higher-quality soft labels introduce significantly fewer mistakes. This aligns with theoretical expectations. More precise soft labels tend to have narrower confidence intervals, which reduces the probability that a sample will flip across the decision threshold due to random annotation variation. As a result, label corrections are more stable, and fewer errors are introduced. We observe even that labeling strategies introduce more new errors than they remove in the original GT for low-quality labels.
@@ VI Conclusion
We introduced RECD, a framework for semi-automated detection and correction of label errors in object detection datasets. For evaluation purposes, we acquired high-quality annotations based on which we identified at least 24.6% missing and inaccurate pedestrian annotations in the original KITTI dataset.
While this work demonstrated the applicability of our framework to detect and correct label errors, we also note that, under loose restrictions, up to 66% of the existing label errors were not found. This limitation stems from the distribution shift between original and validated ground truth, i.e., the predictions are biased by the original ground truth which does not contain small objects in our case. One of the key motivations for releasing our soft-labeled VGT benchmark is to spur the development of better label error detection methods that overcome this recall bottleneck. We considered only a few combinations of object detectors and label error detection methods on a single dataset, focusing on the pedestrian class as an initial use case as it represents a known challenging task. We focused on providing a detailed and focused analysis rather than giving a broad overview. Our focus is on high-quality annotations which is costly and involved plenty of manual / human effort. This cost increases with the number of data points and classes.
Our framework RECD with its modular structure can be used with any object detector network and label error detection method. We demonstrated that it identifies hundreds of label errors faster than humans can annotate bounding boxes. We analyze different annotation procedures regarding label quality and annotation costs and show that attempting to correct datasets with poor-quality annotations may introduce more label errors than it is fixing. We announce the public release of the validated annotations for the pedestrian class for KITTI which includes at least 24.6% missing labels in the original GT. Furthermore, we provide an evaluation platform for the benchmark for label error detection methods for object detection labels introduced in this work.
@@ Acknowledgements
We thank Marius Schubert for discussion and useful advice. S.P. and M.R. acknowledge support by the German Federal Ministry of Education and Research (BMBF) within the junior research group project “UnrEAL” (grant no. 01IS22069). J.K. and M.R. acknowledge support by the German Federal Ministry of Education and Research (BMBF) within the project “RELiABEL” (grant no. 01IS24019B). R.C., D.K. and L.S. also acknowledge support within the project “RELiABEL” (grant no. 01IS24019A).
@@
In Figure A.1 and Figure A.2 you can see examples of the used Annotation interface.
@@
| Object Detector | Accuracy | AUROC |
| Cascade R-CNN | ||
| YOLOX |
-B4 MetaDetect Performance
Meta classification results for the MetaDetect model in terms of Accuracy and AUROC averaged over the folds are listed in Table A.1.
@@
Although not the central goal of this study, we still identified several 85 false positives in the original KITTI dataset based on our validated ground truth (see Figures Figure A.7 and Figure A.8). These findings suggest that our approach can also identify overlabeling errors that would otherwise remain unnoticed.
@@
We analyze cases with low inter-annotator agreement, shown in Figure Figure A.10, and identify common sources of ambiguity in pedestrian detection: distant pedestrians, poor lighting conditions, and crowded scenes with occlusions. Prior work has shown that incorporating soft labels and uncertainty measures can improve model calibration, detection accuracy, and downstream performance [55, 52]. These methods are particularly effective in crowded or visually complex scenes, where binary labels fail to capture annotation uncertainty. Training with soft labels has also been shown to reduce the impact of missing ground truth labels and provide richer training signals by encoding probabilistic information.
@@
Here, we display examples of the label errors we identified by comparing the validated annotations with the original ones. These examples correspond to the values listed in Table 3 of the main manuscript. In the following, we consider only those label errors outside of don’t care regions of the dataset, i.e., the lower half of Table 3. First, we display all of the relatively large overlooked pedestrians and inaccurate annotations for which our validated annotations exhibit a high confidence. Then, we show random examples of other label errors, e.g. representing smaller objects or ambiguous cases, as distinguished in Table 3. We set a common random seed for reproducibility.
-D3 Most evident overlooked objects
In the following, we display the overlooked objects (in red color) that are not contained in original annotations (green color). These examples have a soft label probability of or higher as well as a bounding box height of pixels or more. We identified 63 such cases and order them according to the soft label probability of the validated annotations and include the original filename for verification purposes.
-D4 Most evident misfitting bounding boxes
Original annotations may be inaccurate. If the original annotation does not exhibit an of with the validated one, we consider the original bounding box as misfitting. This label error is less obvious and may depend on subjectivity. We list some of the mismatches we identified, sorted by the difference in with the VGT. Note that many of these errors can be more interpreted as overlooked pedestrians as well but they overlap with an original bounding box annotation mainly due to duplicate GT bounding boxes.
-E Examples of smaller but unambiguous label errors
Here, we consider a random sample of overlooked pedestrians, that were assigned a soft label probability exceeding and have a bounding box height of less than 40 pixels.
-F Examples of somewhat ambiguous label errors
Here, we consider a random sample of overlooked pedestrians, that were assigned a soft label probability between and meaning that they represent somewhat ambiguous instances.

### v2

@@ IV-C Label Error Detection
MetaDetect. We use the MetaDetect meta classification model, which applies gradient boosting. We use the default parameter values and make predictions via -fold cross-validation on the respective hold-out set. The threshold, which determines whether predictions are considered to be TP or FP, is set to 0.5. For each predicted bounding box, the meta classification model predicts a probability for the prediction being correct, based on which we identify label errors. Meta classification results are listed in table III.
| Object Detector | Accuracy | AUROC |
| Cascade R-CNN | ||
| YOLOX |
@@
Instance-wise Loss Method. This method is based on object detection loss at instance level. It requires a two-stage object detector, which is why it is applicable only to Cascade R-CNN predictions. Based on images and the corresponding GT annotations, this method generates label error proposal boxes with a corresponding score and class prediction. We consider proposals with a score of at least 0.01 for the class pedestrian.
@@
ObjectLab. The detected bounding boxes can be evaluated with respect to the overlooked score with ObjectLab. Originally, only confident predictions with a probability greater than 0.95 were considered, but to infer a score for every detected bounding box, this threshold is removed. All other parameters are set to their default values. Additionally, we consider the poor location scores of the original GT annotations. As there is only a single class under consideration, the swapped class label error does not occur here. As label quality scores are normalized and lower values indicate poorer label quality, we subtract these scores from 1 and interpret the result as a probability for the corresponding label issue. The detected bounding boxes, as well as the original GT boxes, are then reviewed in order of probabilities.
IV-D Evaluation Metrics
Label Errors / FN in Original GT: We match predicted bounding boxes with the original GT using an threshold of 0.5, and with the VGT using a threshold of 0.1. These thresholds serve different purposes: the threshold is standard for checking agreement with original annotations, while the threshold allows for approximate matching against VGT, which acts as a proxy for re-annotation. A prediction is considered a label error (i.e., a false negative in the original GT) if it has no match in the original GT but overlaps with the VGT. We classify errors into two types: completely missing () and misaligned boxes (). Matching is performed greedily by descending , and each VGT and GT annotation is matched at most once.
Introduced Label Errors: To assess the risk of new errors introduced by labeling strategies, we compare their outcomes against the VGT. If a semantic validation strategy disagrees with the VGT on a bounding box’s label (e.g., one says pedestrian, the other not pedestrian), the box is misclassified. This may cause errors (FNs or FPs) to be missed or falsely introduced.
V NUMERICAL RESULTS
V-A Comparison of Original and Validated Annotations
Before comparing the VGT to the original GT, we first verify its quality. We evaluate three aspects: whether pedestrians were missed, whether duplicate bounding boxes exist, and, how well the soft labels align with human perception. In a random sample of 200 images, we found that about 0.5% of pedestrians might have been missed, and about 0.8% of pedestrian bounding boxes were duplicates. See the appendix for the full analysis. To evaluate the alignment of soft labels with human perception, we randomly sampled 200 bounding boxes from the VGT and let three expert annotators rate them. Each annotator could choose between the labels ’clearly’, ’maybe’, or ’no’ to indicate whether the bounding box contains a pedestrian. In figure 5, these labels are compared with the corresponding soft label probability. The comparison shows that the categories match well with the expected scores. The category ’maybe’ has a mean and median of 0.48, which is almost perfectly calibrated, as we would expect a value around 0.5. For the category ’clearly’, we would expect a value around 0.8, which is consistent with the results. We conclude that the validated ground truth contains almost no missing pedestrians (about 0.5%), almost no duplicate bounding boxes (about 0.8%), and is well aligned with human perception.
We compare the validated annotations with the original annotations to determine the number of overlooked pedestrians and the number of misfitting or inaccurate bounding boxes. Overlooked pedestrians are VGT annotations that have an of zero with any original annotation. Misfitting boxes are VGT annotations for which there is an intersecting bounding box in the original GT, but with an below 0.5. We distinguish between large and small objects based on the minimum bounding box height, using thresholds of 25 or 40 pixels, which relate to the difficulty levels defined in the KITTI benchmark. Additionally, we consider two different probability thresholds, and , for the soft labels in the VGT. The distribution of soft label probabilities is shown in figure 6(b) and illustrates that while most annotations show high agreement among annotators, some ambiguous cases occur. It is important to note that, unlike the original annotation protocol, the procedure described in IV-B includes all pedestrians, regardless of their size or whether they fall into defined don’t care regions. As a result, we observe that the VGT also contains small objects that were not annotated in the original data, as shown in figure 6(a).
To ensure a fair comparison, we define a default configuration for the annotations under consideration. We include all objects that have an intersection over bounding box area below 0.5 with any don’t care region and a bounding box height of at least 25 pixels. For the VGT annotations, we apply a soft label probability threshold of 0.5 to decide whether to include them. The resulting VGT contains 1,567 annotations for the class pedestrian, which is 671 more than in the original GT. Using this configuration, we identified a total of 384 label errors in the subset of 1,497 KITTI training images (table IV). Out of these, 228 pedestrians were missing entirely and 156 were annotated with inaccurate bounding boxes. We estimate that the original GT misses between 42 and 683 pedestrians, depending on the selected conditions such as minimum object size and probability threshold. This corresponds to between 4 – 76% relative to the 896 original pedestrian annotations. If we treat the overlooked and misaligned annotations as false negatives (FNs) relative to the original GT, we identify 206 FNs under strict conditions (applying a high confidence threshold, excluding don’t care regions, and enforcing size constraints) or 862 FNs when these restrictions are relaxed. This corresponds to a false negative rate (FNs / [FNs + Original GT]) ranging from 18% to 49%, depending on the evaluation setting.
@@ V-A Comparison of Original and Validated Annotations
| Arbitrary | 341 | 173 | 121 | 112 |
| pixels | 228 | 156 | 97 | 109 |
| pixels | 91 | 114 | 42 | 87 |
@@ V-B Label Error Detection and Correction
In figure 7(a), we compare the number of FNs in the original GT detected by various label error detection methods and labeling strategies. As outlined in III, we define label errors as FNs in the original GT. We derive two key insights from the figure: (1) error detection methods make labeling more cost-efficient, and (2) there remains substantial room for improvement.
Among the evaluated methods, only YOLOX+MetaDetect consistently outperforms its base model. Cascade R-CNN+MetaDetect underperforms compared to Cascade R-CNN alone. ObjectLab performs significantly worse, likely due to our adaptation of its method; the original version uses a strict confidence threshold of 0.95, which we relaxed for this analysis. At such high thresholds, models provide highly confident predictions, but they cover fewer objects, limiting error discovery. For instance, using only predictions above 0.95 yields very few detected FNs, even in strong models. This highlights that restricting to high-confidence predictions alone is too limiting, even when using optimal detection techniques. Most detection-based approaches including YOLOX, Cascade R-CNN, and their error detection variants are able to find 150–200 FNs at a cost similar to or lower than the cheapest human annotation strategy. However, the base models saturate at 186 (YOLOX) and 260 (Cascade R-CNN) FNs, while the VGT contains 330 FNs. This indicates that up to 44% of missing labels remain undetected. The gap increases under less restrictive matching conditions (e.g., without minimum box size), where up to 66% of FNs go unnoticed. This limitation stems primarily from the object detectors themselves, which constrain the set of proposed boxes. Labeling strategies show similar trends. As summarized in table I, strategies using keypoint-to-box annotation generate fewer boxes and thus recover fewer FNs. Additional differences can be attributed to variations in label quality across correction strategies (see V-C for ablations). While detectors limit recall, our RECD correction pipeline allows for more precise validation of the boxes they do propose and thus, improves soft labeling quality. In summary, RECD identifies hundreds of label errors in less time than it takes humans to even annotate the bounding boxes. Yet, our analysis also reveals that up to 66% of potentially recoverable FNs remain undetected, highlighting significant research opportunities for improving label error detection.
@@ V-C Trade-offs Between Labeling Strategies and Label Quality
We evaluate the benefits and limitations of various labeling strategies in figure 7(b). As expected, strategies cluster based on their bounding box annotation method discussed in IV-B. This is intuitive, since the number and quality of available boxes heavily influence the ability to identify FNs. Strategies using higher-quality soft labels consistently find more FNs. However, in less restrictive settings, e.g., relaxing the minimum box size, even low-cost strategies can sometimes perform best. It can be observed that the number of newly introduced errors decreases with increasing label quality. The lowest-cost semantic strategy (’Is pedestrian?’) introduces several hundred errors. In contrast, higher-quality soft labels introduce significantly fewer mistakes. This aligns with theoretical expectations. More precise soft labels tend to have narrower confidence intervals, which reduces the probability that a sample will cross the decision threshold due to random annotation variation. As a result, label corrections are more stable, and fewer errors are introduced. We find that low-quality labeling strategies introduce more errors than they remove.
@@ VI Conclusion
We introduced RECD, a framework for semi-automated detection and correction of label errors in object detection datasets. For evaluation purposes, we acquired high-quality annotations based on which we identified at least 18% missing and inaccurate pedestrian annotations in the original KITTI dataset.
This work demonstrated the applicability of our framework to detect and correct label errors. However, we note that up to 66% of the existing label errors remain undetected. This highlights research opportunities for improving label error detection methods. Importantly, RECD is detector-agnostic by design and thus, advances in detection architectures can be leveraged to enhance label error detection. We analyze different annotation procedures with respect to label quality and annotation cost, and show that attempting to correct datasets with poor-quality annotations may introduce more label errors than it fixes.
Future work may extend RECD to additional datasets and classes to assess its transferability, which lies beyond the scope of this work. We announce the public release of the validated KITTI annotations for the pedestrian class. Based on these annotations, we provide a benchmark evaluation platform for label error detection methods for object detection datasets.
@@ Acknowledgements
S.P. and M.R. thank Marius Schubert for discussion and help with baseline implementations. S.P. and M.R. acknowledge support by the German Federal Ministry of Research, Technology and Space (BMFTR) within the junior research group project “UnrEAL” (grant no. 16IS22069). J.K. and M.R. acknowledge support by the BMFTR within the project “RELiABEL” (grant no. 16IS24019B). R.C., D.K. and L.S. also acknowledge support within the project “RELiABEL” (grant no. 16IS24019A).
@@
In figure A.1 you can see an example of the used Annotation interface.
@@
Although not the central goal of this study, we still identified several 85 false positives in the original KITTI dataset based on our validated ground truth (see Figures A.7 and A.8). These findings suggest that our approach can also identify overlabeling errors that would otherwise remain unnoticed.
@@
We analyze cases with low inter-annotator agreement, shown in Figure A.9, and identify common sources of ambiguity in pedestrian detection: distant pedestrians, poor lighting conditions, and crowded scenes with occlusions. Prior work has shown that incorporating soft labels and uncertainty measures can improve model calibration, detection accuracy, and downstream performance [55, 52]. These methods are particularly effective in crowded or visually complex scenes, where binary labels fail to capture annotation uncertainty. Training with soft labels has also been shown to reduce the impact of missing ground truth labels and provide richer training signals by encoding probabilistic information.
@@
Here, we display examples of the label errors we identified by comparing the validated annotations with the original ones. These examples correspond to the values listed in table IV of the main manuscript. In the following, we consider only those label errors outside of don’t care regions of the dataset, i.e., the lower half of table IV. First, we display all of the relatively large overlooked pedestrians and inaccurate annotations for which our validated annotations exhibit a high confidence. Then, we show random examples of other label errors, e.g.representing smaller objects or ambiguous cases, as distinguished in table IV. We set a common random seed for reproducibility.
Most evident overlooked objects. In the following, we display the overlooked objects (in red color) that are not contained in original annotations (green color). These examples have a soft label probability of or higher as well as a bounding box height of pixels or more. We identified such cases and order them according to the soft label probability of the validated annotations and include the original filename for verification purposes.
Most evident misfitting bounding boxes. Original annotations may be inaccurate. If the original annotation does not exhibit an of with the validated one, we consider the original bounding box as misfitting. This label error is less obvious and may depend on subjectivity. We list the first of the most evident mismatches (according to minimal bounding box height and soft label probability) we identified, sorted by the difference in with the VGT. Note that many of these errors can be interpreted as overlooked pedestrians but they overlap with an original bounding box annotation, e.g., due to duplicate GT bounding boxes or overlooked pedestrians being close to annotated ones.
Examples of smaller but unambiguous label errors. Here, we consider a random sample of overlooked pedestrians, that were assigned a soft label probability exceeding and have a bounding box height of less than 40 pixels.
Examples of somewhat ambiguous label errors. Here, we consider a random sample of overlooked pedestrians, that were assigned a soft label probability between and meaning that they represent somewhat ambiguous instances.